OpenStudy (rositaluv):

When the quadratic equation y = (x − 3)(2x + 7) is written in standard form, what is the value of the coefficient "b"?

OpenStudy (anonymous):

standard form,i sy=ax^2+bx+c y=2x^2-6x+7x-21 y=2x^2+x-21 b=1

Parth (parthkohli):

First, you must simplify \((x - 3)(2x + 7)\) \( \color{Black}{\Rightarrow x(2x + 7) - 3(2x + 7) }\) \( \color{Black}{\Rightarrow 2x^2 + 7x - 6x - 21}\) \( \color{Black}{\Rightarrow 2x^2 + x - 21}\)

Parth (parthkohli):

\(b\) is the coefficient of \(x\).

Parth (parthkohli):

What do you find the coefficient of \(x\) to be?

OpenStudy (rositaluv):

1 Thanks guys

Parth (parthkohli):

Glad to help you!

OpenStudy (rositaluv):

I might have more I need help on.